# Towards Robust Automated Perceptual Voice Quality Assessment with Speech Foundation Models

### v2

@@ 4.3 Feature Extraction
All signals were resampled to 16 kHz before feature extraction. Two types of features were extracted. First, SFM embeddings from a pre-trained model (WavLM or Whisper) were used to capture phonetic and prosodic information. These embeddings were computed for both PVQD-A and PVQD-S to evaluate their effectiveness across different speech units. Second, handcrafted acoustic features Jitter, Shimmer, and HNR were extracted using the Praat toolkit [34] through the Parselmouth interface [33],[35].
@@ 5.1 Comparison of VOQANet with the Baseline
Table 5 shows the utterance-level evaluation results of VOQANet with different features (including the embeddings of the last layer of WavLM or Whisper) on CAPE-V and GRBAS prediction. We compare VOQANet with the methods of Lin et al. [16], which include traditional or neural regressors, using handcrafted features (HF), traditional signal transformations (mel-frequency cepstral coefficient and modulation spectrum (MFCC+MS)), raw waveform, and features extracted by pre-trained models (e.g., Wav2Vec2 (W2V2), HuBERT, WavLM, and Whisper). Lin et al. only evaluated their methods on CAPE-V prediction on the PVQD-A subset. Their results show that traditional features (including HF and MFCC+MS) outperform SFM features. SFM features performed worse than expected, except for the representation of the last layer of Whisper (Whisper (Last)). The lowest RMSE is 14.76 and the highest PCC is 0.69. However, their best results are significantly worse than those of VOQANet, suggesting that these traditional features are not sufficient to model the complex acoustic patterns relevant to perceptual voice assessment. Moreover, their models were trained using data augmentation with noise injection, whereas our VOQANet models in this experiment was trained on the original PVQD-A dataset without data augmentation. Despite this difference, VOQANet still performs significantly better than these baselines. On PVQD-A, VOQANet reduces RMSE to 10.514 (Whisper (Last)) and 9.955 (WavLM (WS)), and the latter is more than 30% lower than the best baseline. PCC is also improved substantially, exceeding 0.8 for both embeddings.
@@ 5.2 Comparison of SFM Types and Representations
In this study, WavLM and Whisper are used as backbone SFMs because of their excellent performance in speech quality assessment tasks [30]. Another study also highlighted the growing role of self-supervised learning models in non-intrusive speech assessment [37]. To further analyze the effectiveness of SFM-based representations, we compare two adaptation strategies: last-layer features (Last) and weighted-sum aggregation (WS) in Whisper and WavLM. This experiment adopts utterance-level evaluation. As shown in Table 6, the weighted-sum representation consistently outperforms the last-layer representation in all configurations. For example, for GRBAS prediction on the PVQD-A subset, WavLM WS yields lower RMSE (0.369 vs. 0.380) and higher PCC (0.809 vs. 0.795) than its last-layer counterpart. This is consistent with previous findings in the speech processing literature [17], [18], which show that intermediate transformer layers encode diverse and complementary representations, some of which better capture prosodic and phonatory features critical to the perception of voice quality.
@@ 5.3 Comparison between VOQANet and VOQANet+
While SFM (e.g., WavLM) embeddings are effective in capturing high-level acoustic and prosodic information, they are not explicitly optimized for clinically salient voice quality traits. To address this limitation, VOQANet+ incorporates handcrafted acoustic features (jitter, shimmer, and HNR, or JSH for short). These features have been shown to correlate with perceptual voice quality dimensions and are widely adopted in clinical voice analysis and speech processing systems due to their ability to reflect phonatory stability and noise characteristics [38]. Combined with WavLM embeddings, these handcrafted features provide interpretable low-level signal-based information that complements deep learning-based representations.
As shown in Table 7, VOQANet+ consistently achieves the lowest RMSE and highest PCC across all tasks and evaluation levels, outperforming VOQANet with SFM alone. The performance improvement is particularly prominent at the patient level, where predictions are aggregated over utterances from the same speaker to mimic real-world assessment. For example, in GRBAS prediction on PVQD-S, VOQANet+ improves the patient-level PCC from 0.867 to 0.874 and reduces RMSE from 0.297 to 0.289. These results indicate that adding handcrafted features helps to supplement low-level acoustic cues, such as irregularities in frequency or amplitude, that may not be adequately captured by SFM embeddings alone.
Fig. 3 visualizes the CAPE-V scores predicted by VOQANet versus the actual scores, while Fig. 4 does so for GRBAS. Fig. 5 visualizes the CAPE-V scores predicted by VOQANet+ versus the actual scores, while Fig. 6 does so for GRBAS. In each figure, the x-axis denotes the ground-truth scores assigned by expert raters, and the y-axis represents the model’s predicted scores. The top row (e.g., Fig. 3(a–b)) corresponds to utterance-level predictions, and the bottom row (e.g., Fig. 3(c–d)) shows patient-level predictions obtained by averaging the prediction scores of different utterances of the same speaker. Each point represents an utterance (or speaker), the red line shows the regression fit, and the shaded area is the 95% confidence interval. VOQANet+ shows tighter clustering near the diagonal, especially in the patient-level plots (e.g., Fig. 3(d) vs. Fig. 5(d) and Fig. 4(d) vs. Fig. 6(d)), which indicates stronger prediction consistency and lower variance. These findings further demonstrate that handcrafted features can enhance the model’s ability to estimate perceptual ratings, especially in cases of higher severity, where expert judgments tend to be more variable. In summary, by combining black-box deep learning-based representations with interpretable acoustic features, VOQANet+ improves the clinical relevance, trustworthiness, and robustness of voice quality prediction.
@@
Moreover, the advantage of sentence-based predictions is further amplified when handcrafted features are incorporated into the model (see the comparison of VOQANet and VOQANet+ in Table 7). This suggests a synergistic effect between rich SFM-derived representations and domain-informed acoustic markers, especially in the capture of the characteristics of voice disorders. These findings also have important implications for robustness. As we will explore in the next subsection, sentence-level input enables VOQANet+ to maintain stronger performance under seen and unseen noisy conditions. This resilience further highlights the clinical value of incorporating continuous speech into the automated voice assessment framework.
@@ 5.5 Robustness to Seen and Unseen Noise
This improvement highlights that handcrafted features (JSH), rooted in perturbation measures and periodicity detection, provide complementary information that is robust to noise and remains stable even when deep SFM embeddings degrade. By capturing signal-level voice irregularities that are less susceptible to spectral masking or background interference, VOQANet+ benefits from richer and more noise-resistant representations. Moreover, VOQANet+ shows less performance degradation from seen to unseen noise than VOQANet, especially in GRBAS prediction. This highlights the advantage of combining interpretable clinically meaningful features with learned embeddings to form a more adaptable model that maintains prediction quality even when faces with out-of-distribution acoustic environments. The results confirm that VOQANet+ provides a more robust solution for deployment in noisy real-world environments.
@@ 6 Conclusions
This study introduces VOQANet, a deep learning-based framework with an attention mechanism for automated perceptual voice quality assessment. VOQANet uses SFM to extract representations from speech input, effectively capturing high-level acoustic and prosodic information from raw waveforms. It achieves strong predictive performance on both CAPE-V and GRBAS rating scales, demonstrating the utility of SFM embeddings in modeling perceptual voice characteristics.
To further enhance clinical relevance and model robustness, we propose VOQANet+, which incorporates handcrafted acoustic features (jitter, shimmer, and HNR) and SFM embeddings. Through a comprehensive evaluation on the PVQD dataset, we identify several key findings: (1) WavLM outperforms Whisper in voice quality prediction, especially when using weighted-sum (WS) embeddings that aggregate across multiple layers; (2) VOQANet+ consistently improves performance over VOQANet in both utterance-level and patient-level evaluations, reflecting its better stability and alignment with clinical assessments; and (3) VOQANet+ demonstrates greater resilience under seen and unseen noise conditions, showing greater robustness and generalizability in acoustically challenging environments.
By combining the strengths of pre-trained SFM representations and clinically interpretable acoustic features, VOQANet+ provides a robust and interpretable foundation for real-world voice quality assessment applications. Future research directions may include exploring multi-task learning to jointly predict individual CAPE-V (or GRBAS) dimensions or perceptual subscales, as well as cross-lingual generalization and domain adaptation to enable broader deployment across clinical settings and languages.

### v3

@@ 4.3 Feature Extraction
All signals were resampled to 16 kHz before feature extraction. Two types of features were extracted. First, SFM embeddings from a pre-trained model (WavLM or Whisper) were used to capture phonetic and prosodic information. These embeddings were computed for both PVQD-A and PVQD-S to evaluate their effectiveness across different speech units. Second, LLDs, namely Jitter, Shimmer, and HNR were extracted using the Praat toolkit [35] through the Parselmouth interface [34],[36].
@@ 5.1 Comparison of VOQANet with the Baseline
Table 5 shows the utterance-level evaluation results of VOQANet with different features (including the embeddings of the last layer of WavLM or Whisper) on CAPE-V and GRBAS prediction. We compare VOQANet with the methods of Lin et al. [17], which include traditional or neural regressors, using LLDs, traditional signal transformations (mel-frequency cepstral coefficient and modulation spectrum (MFCC+MS)), raw waveform, and features extracted by pre-trained models (e.g., Wav2Vec2 (W2V2), HuBERT, WavLM, and Whisper). Lin et al. only evaluated their methods on CAPE-V prediction on the PVQD-A subset. Their results show that traditional features (including HF and MFCC+MS) outperform SFM features. SFM features performed worse than expected, except for the representation of the last layer of Whisper (Whisper (Last)). The lowest RMSE is 14.76 and the highest PCC is 0.69. However, their best results are significantly worse than those of VOQANet, suggesting that these traditional features are not sufficient to model the complex acoustic patterns relevant to pathological voice assessment. Moreover, their models were trained using data augmentation with noise injection, whereas our VOQANet models in this experiment was trained on the original PVQD-A dataset without data augmentation. Despite this difference, VOQANet still performs significantly better than these baselines. On PVQD-A, VOQANet reduces RMSE to 10.514 (Whisper (Last)) and 9.955 (WavLM (WS)), and the latter is more than 30% lower than the best baseline. PCC is also improved substantially, exceeding 0.8 for both embeddings.
@@ 5.2 Comparison of SFM Types and Representations
In this study, WavLM and Whisper are used as backbone SFMs because of their excellent performance in speech quality assessment tasks [31]. Another study also highlighted the growing role of self-supervised learning models in non-intrusive speech assessment [38]. To further analyze the effectiveness of SFM-based representations, we compare two adaptation strategies: last-layer features (Last) and weighted-sum aggregation (WS) in Whisper and WavLM. This experiment adopts utterance-level evaluation. As shown in Table 6, the weighted-sum representation consistently outperforms the last-layer representation in all configurations. For example, for GRBAS prediction on the PVQD-A subset, WavLM WS yields lower RMSE (0.369 vs. 0.380) and higher PCC (0.809 vs. 0.795) than its last-layer counterpart. This is consistent with previous findings in the speech processing literature [18], [19], which show that intermediate transformer layers encode diverse and complementary representations, some of which better capture prosodic and phonatory features critical to the perception of pathological voice quality.
@@ 5.3 Comparison between VOQANet and VOQANet+
While SFM (e.g., WavLM) embeddings are effective in capturing high-level acoustic and prosodic information, they are not explicitly optimized for clinically salient voice quality traits. To address this limitation, VOQANet+ incorporates LLDs (jitter, shimmer, and HNR, or JSH for short). These features have been shown to correlate with pathological voice quality dimensions and are widely adopted in clinical voice analysis and speech processing systems due to their ability to reflect phonatory stability and noise characteristics [39]. Combined with WavLM embeddings, these LLDs provide interpretable low-level signal-based information that complements deep learning-based representations.
As shown in Table 7, VOQANet+ consistently achieves the lowest RMSE and highest PCC across all tasks and evaluation levels, outperforming VOQANet with SFM alone. The performance improvement is particularly prominent at the patient level, where predictions are aggregated over utterances from the same speaker to mimic real-world assessment. For example, in GRBAS prediction on PVQD-S, VOQANet+ improves the patient-level PCC from 0.867 to 0.874 and reduces RMSE from 0.297 to 0.289. These results indicate that adding LLDs helps to supplement low-level acoustic cues, such as irregularities in frequency or amplitude, that may not be adequately captured by SFM embeddings alone.
Fig. 3 visualizes the CAPE-V scores predicted by VOQANet versus the actual scores, while Fig. 4 does so for GRBAS. Fig. 5 visualizes the CAPE-V scores predicted by VOQANet+ versus the actual scores, while Fig. 6 does so for GRBAS. In each figure, the x-axis denotes the ground-truth scores assigned by expert raters, and the y-axis represents the model’s predicted scores. The top row (e.g., Fig. 3(a–b)) corresponds to utterance-level predictions, and the bottom row (e.g., Fig. 3(c–d)) shows patient-level predictions obtained by averaging the prediction scores of different utterances of the same speaker. Each point represents an utterance (or speaker), the red line shows the regression fit, and the shaded area is the 95% confidence interval. VOQANet+ shows tighter clustering near the diagonal, especially in the patient-level plots (e.g., Fig. 3(d) vs. Fig. 5(d) and Fig. 4(d) vs. Fig. 6(d)), which indicates stronger prediction consistency and lower variance. These findings further demonstrate that LLDs can enhance the model’s ability to estimate perceptual ratings, especially in cases of higher severity, where expert judgments tend to be more variable. In summary, by combining black-box deep learning-based representations with interpretable acoustic features, VOQANet+ improves the clinical relevance, trustworthiness, and robustness of pathological voice quality prediction.
@@
Moreover, the advantage of sentence-based predictions is further amplified when LLDs are incorporated into the model (see the comparison of VOQANet and VOQANet+ in Table 7). This suggests a synergistic effect between rich SFM-derived representations and domain-informed acoustic markers, especially in the capture of the characteristics of voice disorders. These findings also have important implications for robustness. As we will explore in the next subsection, sentence-level input enables VOQANet+ to maintain stronger performance under seen and unseen noisy conditions. This resilience further highlights the clinical value of incorporating continuous speech into the automated voice assessment framework.
@@ 5.5 Robustness to Seen and Unseen Noise
This improvement highlights that LLD (JSH), rooted in perturbation measures and periodicity detection, provides complementary information that is robust to noise and remains stable even when deep SFM embeddings degrade. By capturing signal-level voice irregularities that are less susceptible to spectral masking or background interference, VOQANet+ benefits from richer and more noise-resistant representations. Moreover, VOQANet+ shows less performance degradation from seen to unseen noise than VOQANet, especially in GRBAS prediction. This highlights the advantage of combining interpretable clinically meaningful features with learned embeddings to form a more adaptable model that maintains prediction quality even when faces with out-of-distribution acoustic environments. The results confirm that VOQANet+ provides a more robust solution for deployment in noisy real-world environments.
@@ 6 Conclusions
This study introduces VOQANet, a deep learning-based framework with an attention mechanism for automated pathological voice quality assessment. VOQANet uses SFM to extract representations from speech input, effectively capturing high-level acoustic and prosodic information from raw waveforms. It achieves strong predictive performance on both CAPE-V and GRBAS rating scales, demonstrating the utility of SFM embeddings in modeling perceptual voice characteristics of disordered voice.
To further enhance clinical relevance and model robustness, we propose VOQANet+, which incorporates LLDs (jitter, shimmer, and HNR) and SFM embeddings. Through a comprehensive evaluation on the PVQD dataset, we identify several key findings: (1) WavLM outperforms Whisper in pathological voice quality prediction, especially when using weighted-sum (WS) embeddings that aggregate across multiple layers; (2) VOQANet+ consistently improves performance over VOQANet in both utterance-level and patient-level evaluations, reflecting its better stability and alignment with clinical assessments; and (3) VOQANet+ demonstrates greater resilience under seen and unseen noise conditions, showing greater robustness and generalizability in acoustically challenging environments.
By combining the strengths of pre-trained SFM representations and clinically interpretable acoustic features, VOQANet+ provides a robust and interpretable foundation for real-world pathological voice quality assessment applications. Future research directions may include exploring multi-task learning to jointly predict individual CAPE-V (or GRBAS) dimensions or perceptual subscales, as well as cross-lingual generalization and domain adaptation to enable broader deployment across clinical settings and languages.